lori was 25 years old when her son jamie was born. in how many years will lori be five years less than three times jamie’s age?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39616)

Is Jamie just born and Lori 25 years old TODAY? If so then In 35 more years, the answer to the example problem's question.

Answer by ikleyn(52776)

.

Let x be the number of years when this event will happen.


Then the Jamie's age is x years;   Lori's age is (x+25),

and the equation is

    x + 25 = 3x - 5.


From the equation,

    25 + 5 = 3x - x

      30   = 2x

       x   = 30/2 = 15.


ANSWER.  In 15 years.

Solved.